Prev Next

SW&IT Meeting Minutes 2020-12-14

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20201130
LR20201130 - Bec, Matthieu 80.0
LR20201130 - Holmberg, Carl 32.0
LR20201130 - Edmeades, David 80.0
LR20201130 - Sola, Igor 40.0
LR20201130 - Grenz, Paul 0.0
LR20201130 - Hooper, Stephen 0.0
LR20201130 - Bryant, Trey 40.0
LR20201130 - Zhang, Yang 80.0
LR20201130 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT SW / ATT PM
    • leadership workshop Tue/Wed 8am-12pm
    • Carl's responsibilities within the group
  • This week
    • ATT SW final edits
    • AHS for TMS, pending PRs
    • iLocater fiber discussions?

PSG

  • Last week
    • Finished up simple terminal program to connect, command, disconnect from d4x5 controller.
  • This week
    • Scheduling time to test program, performing tests and debugging.

DE

  • User accounts
  • UPS batteries
  • Cisco training
This week:
  • Got replacement UPS batteries, will finish packs
  • Office trip this week/early next
  • Cisco training

YZ

  • Previous:
    • Misc TMS work (software updates, wiki edits, meeting...).
    • AO-Cravat/OSCO web UI prototyping.
    • Some LBC work (IT#7595 and RB_Science ongoing updates).
    • Minor TCS change (disabling IIF ICE logging to syslog).
    • Completed UA information security awareness certification.
  • This week:
    • Continue AO-Cravat/OSCO web UI prototyping (and last meeting of the year).
    • Possible TMS software support (and last meeting of the year).
    • Misc LBC and MODS work.

XZ

  • Last week
    • ATT AdSec
    • ATT SW
    • AO operation support
    • UA Information Security Awareness traning
    • Compile blind deconvolution
  • This week
    • ATT AdSec
    • Blind deconvolution test
    • AdSec safety SW changes and test
    • SPIE meeting this week
    • From this Friday, in vacation, holidays, university closure, back to "office" in Jan 11

IS

Last week:
  • LUCI
    • IT 8220- Request to Include Sub-Framing Options in LUCI Software
    • Geirs script.
  • PIT
    • Work on http connection testing
    • Admin tool
  • OT
    • Working on a internal test release for ADI related bug and new AO patrol field

This week:
  • OT
    • Issue new release for ADI fix.
  • LUCI
    • RPM
    • IT 8231 GEIRS timeout
    • IT 8220
    • split current luci motor privilege into two to allow more control on what people can and can't do.
    • Look into reading WebIO status and passing it to the database/telemetry.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
  • PIT
    • pit admin tool
Topic revision: r7 - 14 Dec 2020, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback